OpenStudy (anonymous):

A sample of gas occupies 7.80 liters at 425°C? What will be the volume of the gas at 35°C if the pressure does not change?

OpenStudy (jonaho):

for this you should use the formula V1/T1= V2/T2 convert Celsius to Kelvin 7.8/ 698=V2/ 308 V2=3.44 Liters